Calculate the terminal voltage for a battery with an internal resistance of 0.750 Ω and an emf of 7.00 V when the battery is connected in series with a 71.0-Ω resistor.
Express your answer to three significant figures and include the appropriate units.
Answer #1

Given data

External resistance R = 71.0 ohm

Internal resistance of battery r = 0.750 ohm

Emf of battery E = 7.00 V

Net resistance in series of circuit Ŕ = R + r = 71.0 + 0.750 = 71.750 = 71.8 ohm

Current in the circuit I = E/Ŕ = 7.00/71.8 = 0.0975 A

Therefore

Terminal voltage across battery V = E - Ir = 7.00 - (0.0975)(0.750) = 7.00 - 0.0731 = 6.93 Volt ( rounded off to 3 significant digits)